OpenStudy (anonymous):

Earth Science - After three half-lives, a sample contains 20 grams of parent isotopes. How many grams of the parent isotope were present to start? ANSWERS: 320 160 80 40

OpenStudy (anonymous):

Let there be 'x' grams of isotopes at the beggining. No. of grams remaining after first half life=1/2*x No. of grams remaining after second half life=1/4*x No. of grams remaining after third half life=1/8*x Now,1/8*x=20(given) =>x=160grams